A card is drawn from an ordinary deck of 52 cards, and the result is recorded on paper. The card is then returned to the deck and another card is drawn and recorded. Find the probability that the following occurs.
Only the first card is a club

Answer by stanbon(26259)

A card is drawn from an ordinary deck of 52 cards, and the result is recorded on paper. The card is then returned to the deck and another card is drawn and recorded. Find the probability that the following occurs.
Only the first card is a club
-------------------------------------
Since the card is returned to the deck the events are independent.
-----------------------------------
P(1st card is a club) = 1/4
--------------
P(2nd card is not a club) = 3/4
---------------
P(club and not club) = (1/4)(3/4) = 3/16
==============================================
Cheers,
Stan H.